# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v2

@@ 6.3. Prediction Result
To address RQ3, we proposed several regression models in the experiment. First, a general prediction model was constructed by considering all participants in the study. Second, we built individual prediction models for each participant to account for the unique characteristics and patterns exhibited by different individuals. It is worth mentioning that only mobile phone data was utilized in this study. After excluding 19 records with invalid ESM data, a total of 1456 app-usage data records were obtained at 90-minute intervals for analysis.
@@ 6.3.1. General Model
The prediction performance for the general models is presented in Figure 8. The models used in the comparison include Mean baseline, Median baseline, KNN regressor, RF regressor, GB regressor. Among these models, the RF regressor demonstrated superior performance to both baselines in terms of MAE and RMSE for predicting valance and arousal. Interestingly, the RF regressors with different feature sets achieved comparable performance, with all feature sets yielding the best results. This can be attributed to the capability of RF to assign weights to features based on their importance, where incorporating more features leads to improved prediction performance. Furthermore, it is observed that the proposed model generally exhibited better performance in predicting valence compared to arousal, the difference in performance may be attributed to the complexity of arousal, which is influenced by various factors such as physiological arousal and cognitive processing. These factors may not be fully captured by mobile usage behaviours alone.
@@
Table 5 displays the prediction results of the general models for valence and arousal. Regarding the valence dimension, our results indicate that the RF regressor achieved the best predictive performance across three different feature sets. For the regression model using all features, the RMSE obtained by the RF algorithm was 0.903, surpassing the Median baseline by 17.23%. Furthermore, the MAE of the RF regressor using all feature sets was 0.717, outperforming the mean baseline by 12.99% and the median baseline by 6.52%. Similar results were observed for the models utilizing the top 17 features. Regarding the arousal dimension, our results indicate that the RF regressor utilizing all features achieved the best performance in terms of RMSE, with a value of 1.005. This outperformed the Median baseline by 12.30%. These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
@@ 6.3.2. Individual Models
The individual models for both valence and arousal were trained using all 29 features. However, due to some participants having insufficient data recorded, the models were only trained on those with more than 50 data points. This resulted in a total of 13 participants and 1220 data points being used. The prediction results for these models, with the ID of participants indicated on the y-axis, can be seen in Figure 9. Overall, we can observe variations in the prediction performance of valence and arousal across different participants. For instance, participant ‘M11OAR’ consistently exhibits higher MAE and RMSE on both valence and arousal than the other participants, e.g., ‘H12DPB’ and ‘E02OKG’. This variability suggests that individual differences play a significant role in determining the accuracy of the predictions, highlighting the need to consider personalized factors when developing prediction models for emotion in home confinement situations.
@@
(b) MAE for arousal
(c) RMSE for valence
@@ 6.4. Feature Importance
To interpret the key factors influencing the prediction of valence and arousal during home confinement, we examine the feature importance of the best-performing models, which in this case are the RF regressors. For brevity, we focus on the top 15 features for each model, as shown in Figure 10. The RF regressors utilize impurity-based feature importance, which measures how much including a particular feature reduces impurity across the decision tree nodes. A higher score indicates a more important feature, as it leads to a greater reduction in impurity. A higher score indicates a more important feature, as it leads to a greater reduction in impurity. As depicted in Figure 10, we observe that the most important features for both valence and arousal are similar, but with different weights. These features primarily include time-related features, app-category features, and notification features. These findings suggest that these variables play a significant role in accurately predicting participants’ valence and arousal levels during home confinement.
(a) Valence
(b) Arousal
@@
Figure 10. Top 15 features based on impurity score with RF regressors
@@
Both models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day significantly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less significance in the arousal model. Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features, suggesting that while there is still variation in terms of time and day, weekdays may not have a significant impact on participants’ valence and arousal levels. These findings differ slightly from the observations reported in (44), which found that app usage patterns varied greatly by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in (18), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
The above findings have important implications for understanding the factors that contribute to human well-being, particularly in the context of home confinement. They provide valuable insights that can inform the development of interventions aimed at promoting emotional wellness during such periods. By considering the impact of app usage, notifications, and time-related factors, we can also design targeted strategies to enhance individuals’ emotional well-being and overall quality of life.
7. Implications
Our study examines the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance. Through the use of unobtrusive mobile phone sensing, we aim to predict participants’ well-being and social roles. First of all, our findings highlight the importance of mental health during periods of home confinement. Home confinement is not limited to temporary pandemic adaptation but also the growing trend of working from home post-pandemic (40; 69) and other similar situations, such as post-operative recovery, parental leave, or limited mobility related to disability, or even mobile phone addiction. Secondly, our quantitative approach and regression models can contribute to tracking, early detection and treatment of mental illnesses during these circumstances. Our results also help policymakers to develop appropriate strategies that minimize the long-term negative effects of the pandemic on mental health and encourage people to build healthy digital habits (69). For example, quarantine or isolation time should be kept as short as possible, and voluntary quarantine could be promoted to reduce long-term mental health complications, as recommended by Brooks et al. (12).
In the post-pandemic era, the traditional office-only working arrangement has shifted to more flexible arrangements, with an increasing number of people opting to work from home (53; 40). Studies have shown that individuals who continued to work from home after the pandemic exhibited similar behaviours to those during the pandemic (57), underscoring the importance of studying people’s behaviours in high technology-reliance contexts. Our study was conducted in a city with the longest home confinement period, offering optimal experimental conditions to investigate the influence of app usage on well-being while minimizing confounding effects from other variables, such as nature and physical social interactions. Furthermore, our study contributes to understanding the mental health benefits and consequences of working from home or relying heavily on technology. For instance, tracking mental well-being status provides a means of measuring work-life balance and enables individuals to manage their working schedules or arrangements (e.g., work-from-home or half-office half-home) more effectively.
In addition, this study provides a valuable opportunity for researchers to study the psychological states and smartphone usage behaviours of individuals under home confinement conditions, which has important implications for various fields. For example, it may inform the development of interventions to support individuals in similar situations, such as those willing or undergoing post-operative home rehabilitation. Rethinking about the future of work and life scenarios, where heavily rely on digital technologies. It is important to understand how people deal with being confined to their homes and how technology impacts their lives. This understanding can help researchers design smart environments, or assist healthcare professionals in developing effective remote support programs that enhance the quality of care for those in similar situations. Furthermore, the findings from this study may have broader implications for our understanding of human behaviour and mental health during times of crisis, informing the development of new policies and interventions to support individuals in future crises.
@@ 8. Limitations
We have identified three main limitations in this research. One significant limitation is the selection of time and place. Regarding time, we were unable to compare user well-being and digital device usage data before COVID-19 lockdown with the data during the lockdown due to the unpredictable nature of lockdown policies. However, we have mitigated this limitation by allowing users to report the impact of the lockdown on their well-being and work-life balance through self-reported questionnaires. These data clearly demonstrate the effects of lockdowns on people’s lives, as discussed in Section 4. In terms of geography, our data collection was limited to Melbourne, where the COVID-19 lockdown with Stage IV restrictions was applied. While the specific measures taken by different countries and regions vary, and the overall mental well-being of populations may differ, studies conducted in other countries have also demonstrated the impact of COVID-19 on human well-being and mental health (11; 15; 12). Therefore, our research still holds value in informing government officials or managers to make appropriate decisions and highlighting the importance of mental health care during home confinement situations.
The second limitation is the relatively short duration of the data collection period. We chose to collect data for only three weeks in order to minimize interruptions from ESM questionnaires and end-of-day surveys, with the goal of reducing the burden on participants during the challenging realities of COVID-19 home confinement. Despite the short duration, we were still able to collect a substantial amount of data, including 502,485 records of users’ context and phone usage, 1,749 ESM responses, and 265 end-of-day survey responses. It is worth noting that the delivery of the end-of-day survey link through emails allowed users to complete the survey voluntarily at their convenience. Although the number of end-of-day survey responses is small, the optional nature of the surveys, especially text questions, ensures the credibility and reliability of the collected data. However, due to the limited three-week period of our data collection, it is important to acknowledge that if strict home confinement lasts longer, the emotional experiences of individuals may evolve as they adapt to new challenges.
Finally, the third limitation of this study is the number and representativeness of the participants. A total of 32 participants were recruited, with 27 using smartphones, 13 using desktops, and only 8 using both. Since the majority of participants use mobile phones for data collection, the majority of our background application data analysis focuses on mobile usage data. Additionally, participants were primarily recruited through Facebook and Discord advertisements, which may have led to a population that is skewed towards tech-savvy information workers. Additionally, there is a possibility of oversampling individuals who were ‘relieved to be safe at home and cherish the opportunity to spend more quality time with their family,’ as these individuals may have had more availability to participate in the study. It is important to acknowledge that the COVID-19 lockdown presented challenges in recruiting participants. However, despite these challenges, we were able to collect sufficient data from each participant to conduct our analysis.
@@ 9. Conclusion
In this study, we delved into how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period. Our proposed models yielded promising results in predicting human emotions during home confinement. Moreover, we uncovered intriguing insights regarding the correlation between different factors and human well-being. To address our research objectives on home confinement, we formulated three research questions and key findings are summarized below.
In response to RQ1, we found that participants reported relatively high levels of valence but relatively low levels of arousal, suggesting that participants experienced a state of calmness and relaxation during COVID-19 home confinement. Additionally, our analysis suggested that the COVID-19 lockdown has a detrimental effect on valence and work-life balance, with participants experiencing longer work hours due to interruptions and a blurring of personal and work boundaries. Regarding RQ2, we segmented app usage data and identified the most frequent user activities ‘Just Communication’ and ‘Idling’, potentially attributed to increased virtual communication and available free time. We also found that tools-related activities positively correlated with valence, suggesting people used productivity apps effectively. Moreover, interruptions displayed a positive association with Communication Tools, possibly due to the social nature of these tools leading to more interruptions during work hours. Moving on to RQ3, we developed regression models to predict people’s valence and arousal based on their mobile usage during home confinement. These models exhibited superior performance compared to baseline models. Notably, our findings revealed that time-related features emerged as the primary predictors for both valence and arousal.

### v3

@@ 5.3.1. General Model
After excluding 19 records with invalid ESM data, a total of 1456 app-usage data records were obtained at 90-minute intervals for analysis were used to build the general models.
@@
The prediction performance for the general models is presented in Table 6. The models used in the comparison include Mean baseline, Median baseline, KNN regressor, RF regressor, GB regressor. Among these models, the RF regressor demonstrated superior performance to both baselines in terms of MAE and RMSE for predicting valance and arousal. Interestingly, the RF regressors with different feature sets achieved comparable performance, with all feature sets yielding the best results. This can be attributed to the capability of RF to assign weights to features based on their importance, where incorporating more features leads to improved prediction performance. Furthermore, it is observed that the proposed model generally exhibited better performance in predicting valence compared to arousal, the difference in performance may be attributed to the complexity of arousal, which is influenced by various factors such as physiological arousal and cognitive processing. These factors may not be fully captured by mobile usage behaviours alone.
Regarding the valence dimension, our results indicate that the RF regressor achieved the best predictive performance across three different feature sets. For the regression model using all features, the RMSE obtained by the RF algorithm was 0.903, surpassing the Median baseline by 17.23%. Furthermore, the MAE of the RF regressor using all feature sets was 0.717, outperforming the mean baseline by 12.99% and the median baseline by 6.52%. Similar results were observed for the models utilizing the top 17 features. Regarding the arousal dimension, our results indicate that the RF regressor utilizing all features achieved the best performance in terms of RMSE, with a value of 1.005. This outperformed the Median baseline by 12.30%. These outcomes suggest that the selected features serve as effective predictors of arousal for participants during home confinement.
5.3.2. Group Models
We gathered participant work start times from pre-study questionnaires and categorized participants into two distinct groups: $EARLY$ Workers (678 data points from 12 participants who typically commence work between 7:00 and 10:00 AM) and $LATE$ Workers (778 data points from 13 participants who begin work after 3:00 PM). For the early workers, the average self-reported $valence$ is 3.2 (±$\pm 0.9$) and $arousal$ is 2.9 (±$\pm 1.0$). For the late workers, the average $valence$ is 3.7 (±$\pm 1.0$) and $arousal$ is 2.6 (±$\pm 1.2$). Prediction models were subsequently developed for each group using all features.
As shown in Figure 7(a), when predicting $valence$ for early workers, the models performed similarly to the baselines. However, for late workers, the RF regressor outperformed the Mean baseline, achieving a 13.82% reduction in RMSE and a 20.36% reduction in MAE. A similar pattern is observed in Figure 7(c). For $arousal$ prediction, the models demonstrated limited predictive ability for early workers but performed better for late workers, with the RF regressor reducing RMSE by 17.88% and MAE by 27.74% compared to the Mean baseline. Compared to the general models, there was no improvement in $valence$ prediction for early workers, whereas for late workers, the models showed slight improvements (5% in RMSE) in both $valence$ and $arousal$ predictions.
(a) RMSE for valence
(b) MAE for valence
(c) RMSE for arousal
(d) MAE for arousal
Figure 7. Prediction performance of $valence$ and $arousal$ models for two user groups: $EARLY$ workers (typically starting work between 7:00 and 10:00 AM) and $LATE$ workers (starting work after 3:00 PM). The models are evaluated with 3-fold cross-validation, where the error bar presents the standard deviation.
These results highlight the potential value of considering user groups and routines when studying emotions. Specifically, the findings suggest that early workers’ emotions may be influenced by factors beyond app usage, such as external environmental cues during the day. In contrast, late workers may exhibit a stronger link between app usage behaviours and emotions. Further insights will be explored in the later section (Section 5.4) by analyzing feature importance.
@@ 5.3.3. Individual Models
The individual models for both valence and arousal were trained using all 29 features. Due to insufficient data for some participants, only those with over 50 data points were included, resulting in a total of 13 participants and 1,220 data points. The prediction results for these models, with the ID of participants indicated on the y-axis, can be seen in Figure 8. Overall, we can observe variations in the prediction performance of valence and arousal across different participants. For instance, participant ‘M11OAR’ consistently exhibits higher MAE and RMSE on both valence and arousal than the other participants, e.g., ‘H12DPB’ and ‘E02OKG’. This variability suggests that individual differences play a significant role in determining the accuracy of the predictions, highlighting the need to consider personalized factors when developing prediction models for emotion in home confinement situations.
@@
(b) RMSE for valence
(c) MAE for arousal
@@ 5.4. Feature Importance
To interpret the key factors influencing the prediction of valence and arousal during home confinement, we examine the feature importance of the best-performing models, which in this case are the RF regressors. The RF regressors utilize impurity-based feature importance, which measures how much including a particular feature reduces impurity across the decision tree nodes. A higher score indicates a more important feature, as it leads to a greater reduction in impurity. For brevity, we focus on the top 15 features for the general model and compare them with the top 5 features for the group models.
@@
Figure 9. Top 15 features based on impurity score with RF regressors of the general models.
As depicted in Figure 9, we observe that the most important features for both valence and arousal are similar, but with different weights. These features primarily include time-related, app-category, and notification features. Both general models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day greatly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less importance in the arousal model.
@@
Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features in the general models, but it appears as important in each group models. It worth mentioning that ‘time_period’ is the most important feature for predicting $valence$ for $EARLY$ workers, but not for $LATE$ workers. These findings differ slightly from the observations reported in (35), which found that app usage patterns varied greatly by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in (12), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
Figure 10. Top 5 features based on impurity score with RF regressors of group models.
Besides, as shown in Figure 10, for $LATE$ workers, the same five features with similar importance distributions are observed for both $valence$ and $arousal$ predictions, all are related to app interaction. In contrast, $EARLY$ workers show a more diverse feature set; while the interacted app notification is the second most important, physical activity also contribute to $valence$ prediction. This difference may reflect the broader range of daytime activities available to $EARLY$ workers, potentially explaining their lower prediction performance discussed in Section 5.3.2. For $LATE$ workers, the reliance on app interaction highlights a potential concern regarding phone dependency.
6. Discussion and Implications
This study explores the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance, using unobtrusive mobile phone sensing to gain insights into participants’ emotional states and behaviors. Although the study specifically focuses on the COVID-19 lockdown, we argue that its findings have broader relevance, especially in light of the growing trend of remote work and other situations of enforced isolation.
6.1. Well-being and Social Roles During Home Confinement (RQ1)
Our analysis revealed that participants generally experienced a state of calmness and relaxation, reflected in relatively high levels of valence (positive emotional state) and lower levels of arousal (indicating a lack of intense emotional stimulation). While this suggests a degree of emotional stability during the lockdown, the study also highlights the negative effects of home confinement on work-life balance. Participants reported longer work hours and the blurring of personal and professional boundaries, which resonates with the challenges faced by remote workers in post-pandemic contexts. The segmentation of app usage data also identified key behaviors, such as increased use of communication apps and idle device time, which may reflect both increased virtual socialization and available free time.
Although the findings are rooted in the context of the COVID-19 lockdown, the implications extend beyond this unique situation. Home confinement is not limited to pandemic scenarios but is increasingly relevant to other forms of isolation or reduced mobility, such as remote work, post-operative recovery, disability-related confinement, or even digital addiction. The growing trend of telework and flexible work arrangements in the post-pandemic world further underscores the importance of understanding the emotional and behavioral impacts of prolonged confinement in various forms.
6.2. Predicting Emotional States Through Mobile Usage (RQ2)
In response to RQ2, we developed regression models that predict participants’ valence and arousal based on their mobile usage data. These models demonstrated strong predictive performance, emphasizing the value of mobile phones as a tool for tracking emotional well-being in real time. Our findings indicate that time-related features (e.g., the timing of mobile phone usage) were key predictors of both valence and arousal, followed by app notifications and interactions with those notifications. This temporal sensitivity suggests that personalized interventions based on users’ specific routines could enhance emotional monitoring and support during home confinement.
One key finding was the variability in the performance of prediction models across different user groups. For example, users who start work later in the day relied more heavily on mobile phone interactions, while those with early start times would benefit from models that incorporate data from other activities, such as physical exercise or interactions with environmental cues. This insight is crucial in extending the applicability of our findings to various forms of isolation beyond the pandemic. Whether it’s remote workers, individuals recovering from surgery, or people experiencing limited mobility due to health conditions, the need for individualized and context-aware models remains a central takeaway.
6.3. Design Implications
Based on the findings of this study, we propose several design implications for future interventions and mobile applications aimed at supporting well-being during periods of home confinement or isolation:
Personalized Predictive Models. The success of time-related features in predicting emotional states suggests that mobile apps should incorporate personalized, time-sensitive data to offer real-time emotional tracking and tailored interventions. For example, apps could adjust their notifications and suggestions based on users’ routines, ensuring that interventions are timely and relevant.
Balanced Digital Interactions. Our study found that interruptions, particularly from communication apps, may exacerbate the blurring of work-life boundaries. Future apps should focus on helping users manage interruptions, offering customizable notification settings or “focus mode” features to reduce disruptions and help maintain better separation between work and personal life.
Encouraging Healthy Digital Habits. Tools that support productivity, such as task managers and calendars, were associated with higher valence. Future designs should encourage healthy digital habits by integrating features that promote structure, goal-setting, and task management. Additionally, incorporating reminders for physical activity or mental health breaks could further support users’ emotional well-being during periods of isolation.
Context-Aware Interventions. Given the differences in how users experience isolation based on their routines, digital interventions should be context-aware and adaptive. For example, users with later work hours may benefit from interventions encouraging physical activity or relaxation earlier in the day, while users with earlier start times may need more focused strategies for managing digital distractions during the day.
6.4. Generalizability of the Findings
Although this study specifically focuses on the COVID-19 home confinement context, we believe the findings can extend to a range of other isolation scenarios. While the emotional triggers—such as fear of infection in a pandemic or frustration with limited mobility in recovery—may differ, the core psychological mechanisms and emotional dynamics of isolation are broadly similar. Our study observed patterns like emotional detachment, stress, and blurred work-life boundaries, all of which are commonly experienced during prolonged isolation, regardless of the specific circumstances.
The emotional dynamics of isolation—such as changes in social interactions, work-related stress, and adjusting to new routines—are consistent across scenarios like remote work, post-surgical recovery, or other forms of confinement. For example, both remote workers and those recovering from surgery face challenges with work-life balance and social isolation. Our findings about app usage—especially productivity apps promoting positive emotional states—are relevant in these contexts, suggesting that digital tools can support well-being across various isolation scenarios.
Additionally, the mobile sensing approach used in this study, which tracks behaviors like app usage and interaction timing, can be applied beyond COVID-19. While the types of apps and behaviors may vary, the principle of using mobile devices to track emotional states in real-time is adaptable to different situations, such as remote work or recovery periods. The ability to monitor and intervene based on individual usage patterns could be valuable in a wide range of confinement contexts.
Lastly, context-aware interventions are crucial for emotional well-being during isolation. Our study highlights the importance of personalized, real-time interventions—whether it’s through managing digital interruptions, promoting relaxation, or encouraging physical activity. These strategies can be adapted for any form of isolation, from remote work to recovery, offering broader applicability for improving mental health.
@@ 7. Limitations
This study has several limitations that must be acknowledged, which may affect the validity of our findings (78):
Time and Geographical Constraints (Threats to Construct and External Validity). The timing and location of data collection limit the generalizability of our results. We could not compare users’ well-being and digital device usage before and during the COVID-19 lockdown due to the unpredictability of lockdown policies. Although participants reported their experiences via self-reported questionnaires, the effects of different lockdown measures and their duration may vary across regions. As the study was limited to Melbourne, the findings may not fully reflect experiences in other areas with differing restrictions.
Short Duration of Data Collection (Threats to Internal Validity). The data collection period was limited to three weeks to minimize participant burden during home confinement. While this yielded a substantial dataset, a longer observation period would have provided deeper insights into how digital device usage and emotional well-being evolve over time, particularly as individuals adapt to prolonged lockdowns. Future studies should consider longer data collection windows to capture these potential changes.
Device Usage and Untracked Data (Threats to Internal and Construct Validity). A key limitation is the potential for incomplete data due to participants using untracked devices, such as tablets or additional smartphones, or even desktops while only contributing smartphone data. These unmonitored interactions could have influenced emotional states and introduced noise into the data. The absence of comprehensive tracking across all devices limits our ability to fully assess participants’ total digital behavior and emotional responses, suggesting the need for more extensive device tracking in future studies.
Sample Representativeness (Threats to External and Construct Validity). The sample size (32 participants) and recruitment methods (via Facebook and Discord ads) may have resulted in a skewed population, likely favoring tech-savvy individuals, particularly information workers. Additionally, participants who felt more positively about being at home during the lockdown may have been overrepresented. These factors limit the generalizability of the findings to a broader population. However, despite these challenges, the data collected was sufficient for analysis.
@@ 8. Conclusion
In this field study, we explored how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period. This rare event offers optimal conditions to while minimizing confounding factors like nature and physical social interactions. It contributes to understanding the mental health implications of heavy reliance on technology and work-from-home arrangements. By tracking mental well-being, the study highlights how individuals manage work-life balance and adapt their schedules or arrangements (e.g., work-from-home or half-office half-home) effectively. Additionally, it provides valuable insights into psychological states and smartphone usage behaviours during confinement, informing interventions for similar scenarios, such as post-operative home rehabilitation. Understanding how people navigate home confinement and the role of technology in their lives can guide the design of smart environments and remote support programs, improving care quality. These findings also have broader implications for understanding human behaviour during crises, supporting the development of policies and interventions for future challenges.